# Feedforward equilibrium trajectory optimization with GSPulse

### v1

@@ 3.4 SPARC scenario design: plasma breakdown and startup
Feedforward is most commonly employed during the early phases of a pulse, when the sequencing for plasma breakdown and startup is critical, and at low before reliable equilibrium reconstructions are available. GSPulse is a natural tool for designing the electromagnetic trajectories required for plasma breakdown since this is a trajectory optimization of the vacuum fields [29]. Figures 8 and 9 illustrate an example designing plasma breakdown conditions for SPARC. The electromagnetic conditions required for reliable plasma startup can be summarized as [30, 31]:
These are “instantaneous” conditions that must be satisfied near the plasma formation point (typically near the center of the machine or on the inboard limiter) in order to achieve main ion burn-through and plasma breakdown. However, in practice a number of additional time-dependent characteristics are necessary or desirable. These include: avoiding accidentally creating breakdown conditions (reverse breakdown) before the desired start time, sustaining high electric field for a sufficiently long time to burn through the main plasma ion impurities, achieving a vertical field evolution that provides radial force balance as the plasma current channel forms and grows, achieving a radial field evolution that transitions from field null conditions to providing passive vertical stability on a reasonable timescale such that the growing plasma does not go unstable, and pre-charging and balancing the coil currents to maximize the flux that is available for inductive current drive during the pulse. For additional detail on specifying the breakdown trajectory targets, we refer readers to Leuer’s work for KSTAR and EAST [31]. However, assuming that the target electric and magnetic field evolution have been identified with the appropriate physics modeling, finding the feedforward currents and voltages is easily transformed to a target-tracking optimization problem for GSPulse.
@@
The location for evaluating and could be either area-averaged quantities usually near the center of the tokamak or at specified locations and is generally an operator preference or tuning parameter. The target trajectories for SPARC at specified locations are shown in fig. 9 and indicate that the breakdown conditions can be reliably met on SPARC. These results are generally consistent with previous breakdown simulations where coil current and voltage trajectories were tuned by hand [32, 33]. The peak electric field is 0.9V/m and limited by the power supply voltage capabilities of the CS1 coil. This method is able to find a larger area-averaged than was previously found by hand. However, the main advantage is that the operator will be able to tune the breakdown scenario by specifying the desired signal waveforms directly, instead of having to simultaneously adjust and balance multiple coils.
@@ A.1 Model equations:
3 for second-order difference () terms |
@@ A.2 Algorithm Steps
This step updates the coil currents and applied flux, and is done by casting the update into the form of a quadratic optimization problem similar to model predictive control.
@@
The conductor evolution is written for both coil and vessel elements. Combining eqs. 14 and 15 and being explicit about coils or vessel elements, we arrive at:
@@
Now, lets turn our attention to the dynamics. Using a technique from Model Predictive Control, we can form a prediction model of the dynamic evolution. From the state-space dynamics equation we have that:
@@
Then, plugging in eqs. 36 and 34 into eq. 30, we arrive finally at the standard form of the quadratic program:
@@
Assuming that the state-space output matrix is linear, then stays constant across iterations and only (eq. 35) needs to be updated. This means that many of the matrix objects can be pre-computed only one time and re-used, saving on computational time since a large fraction of the compute is absorbed by forming and multiplying the large matrices. Therefore it is very advantageous to ensure that the output matrix is linear and does not need to be updated, which is an advantage for specifying outputs in terms of flux units instead of spatial unit. For outputs that have units of B-field, flux, or current, then is linear.
@@
After using the optimizer to obtain the updated conductor currents, we compute from eq. 42 which is the partially-updated flux on the grid. Then we use touch-point finder, x-point finder, and o-point finder algorithms on this flux distribution. Out of the identified potential touch-points, x-points, and o-points, we select the magnetic axis and boundary-defining point of the flux distribution (where the boundary-defining point is either a touch-point or an x-point). The algorithm to identify which points are the magnetic axis and boundary follows the logic presented in [18] section 2.3, which is an efficient method for identifying the boundary. The output of all these steps is to identify the flux at the magnetic axis and the flux at the boundary as well as to identify the physical boundary of the plasma.
@@
The linear system of 2 equations (eqs. 50 and 48) and 2 unknowns can be solved to obtain the profile coefficients. This gives us the necessary total and profiles, which are used to obtain the total flux distribution according to eq. 43.

### v2

@@ 3.4 SPARC scenario design: plasma breakdown and startup
Feedforward is most commonly employed during the early phases of a pulse, when the sequencing for plasma breakdown and startup is critical, and at low before reliable equilibrium reconstructions are available. GSPulse is a natural tool for designing the electromagnetic trajectories required for plasma breakdown since this is a trajectory optimization of the vacuum fields. Similar optimization work has also been performed for other tokamaks in di Grazia et al. [2024, 2025], Leuer et al. [2010]. Figures 9 and 10 illustrate an example designing plasma breakdown conditions for SPARC. The electromagnetic conditions required for reliable plasma startup can be summarized as Lloyd et al. [1991], Leuer et al. [2010]:
These are “instantaneous” conditions that must be satisfied near the plasma formation point (typically near the center of the machine or on the inboard limiter) in order to achieve main ion burn-through and plasma breakdown. However, in practice a number of additional time-dependent characteristics are necessary or desirable. These include: avoiding accidentally creating breakdown conditions (reverse breakdown) before the desired start time, sustaining high electric field for a sufficiently long time to burn through the main plasma ion impurities, achieving a vertical field evolution that provides radial force balance as the plasma current channel forms and grows, achieving a radial field evolution that transitions from field null conditions to providing passive vertical stability on a reasonable timescale such that the growing plasma does not go unstable, and pre-charging and balancing the coil currents to maximize the flux that is available for inductive current drive during the pulse. For additional detail on specifying the breakdown trajectory targets, we refer readers to Leuer’s work for KSTAR and EAST Leuer et al. [2010]. However, assuming that the target electric and magnetic field evolution have been identified with the appropriate physics modeling, finding the feedforward currents and voltages is easily transformed to a target-tracking optimization problem for GSPulse.
@@
The location for evaluating and could be either area-averaged quantities usually near the center of the tokamak or at specified locations and is generally an operator preference or tuning parameter. The target trajectories for SPARC at specified locations are shown in fig. 10 and indicate that the breakdown conditions can be reliably met on SPARC. These results are generally consistent with previous breakdown simulations where coil current and voltage trajectories were tuned by hand Wai [2023], Wai et al. [2022]. The peak electric field is 0.9V/m and limited by the power supply voltage capabilities of the CS1 coil. For and , the most critical phase is from ms onwards and we only apply high weights during this period, allowing for some tracking error mostly present in before this time. Before ms, the main criterion is to have nonzero field in order to prevent reverse breakdown conditions Battaglia et al. [2019] where a reverse bias of the electric field can cause premature breakdown. Note that this is not strictly required for this scenario where is approximately zero before breakdown, which is achieved by holding the CS1 current steady at max current, but this is not true for all SPARC breakdown scenarios. GSPulse is able to find a larger area-averaged than was previously found by hand. However, the main advantage is that the operator will be able to tune the breakdown scenario by specifying the desired signal waveforms directly, instead of having to simultaneously adjust and balance multiple coils.
@@ A.1 Model equations:
| 3 for second-order difference () terms | ||||
@@ A.2 Algorithm Steps
This step updates the coil currents and applied flux, and is done by casting the update into the form of a quadratic optimization problem using a technique similar to that used in Model Predictive Control (MPC).
@@
The conductor evolution is written for both coil and vessel elements. Combining eqs. 16 and 17 and being explicit about coils or vessel elements, we arrive at:
@@
Now, lets turn our attention to the dynamics to “unroll” or project forward how future states depend on the sequence of actuator actions. This step is adapted from MPC, where the unrolled system is called the prediction model in MPC contexts. From the state-space dynamics equation we have that:
@@
Then, plugging in eqs. 38 and 36 into eq. 32, we arrive finally at the standard form of the quadratic program:
@@
Assuming that the state-space output matrix is linear, then most terms within the QP problem formulation (eqs. 39, 40, 41 and 42) are constant across iterations. The only nonlinear term that must be updated between iterations is (eq. 37). This means that most terms within the QP can be computed only once and re-used, saving on computational time since a large fraction of compute is absorbed by obtaining the linearization and multiplying large matrices to form the QP. Therefore it is very advantageous to ensure that the output matrix is linear and does not need to be updated. This is the motivation within GSPulse for specifying the error signals in terms of isoflux-type outputs (flux and field differences) instead of spatial gap-related units. For outputs that have units of B-field, flux, or current, then is linear because of the linearity of the Green’s functions.
@@
After using the optimizer to obtain the updated conductor currents, we compute from eq. 44 which is the partially-updated flux on the grid. Then we use touch-point finder, x-point finder, and o-point finder algorithms on this flux distribution. Out of the identified potential touch-points, x-points, and o-points, we select the magnetic axis and boundary-defining point of the flux distribution (where the boundary-defining point is either a touch-point or an x-point). The algorithm to identify which points are the magnetic axis and boundary follows the logic presented in Moret et al. [2015] section 2.3, which is an efficient method for identifying the boundary. The output of all these steps is to identify the flux at the magnetic axis and the flux at the boundary as well as to identify the physical boundary of the plasma.
@@
The linear system of 2 equations (eqs. 52 and 50) and 2 unknowns can be solved to obtain the profile coefficients. This gives us the necessary total and profiles, which are used to obtain the total flux distribution according to eq. 45.